Determining Manufacturing Costs

Wagner Printing Inc. began printing operations on July 1. Jobs 7-01 and 7-02 were completed during the month, and all costs applicable to them were recorded on the related cost sheets. Jobs 7-03 and 7-04 are still in process at the end of the month, and all applicable costs except factory overhead have been recorded on the related cost sheets. In addition to the materials and labor charged directly to the jobs, $1,200 of indirect materials and $4,300 of indirect labor were used during the month. The cost sheets, in summary form, for the four jobs during the month are as follows:

Job 7-01 Job 7-02
Direct materials 12,400 Direct materials 5,800
Direct labor 4,800 Direct labor 2,500
Factory overhead 2,640 Factory overhead 1,375
Total 19,840 Total 9,675
Job 7-03 Job 7-04
Direct materials 15,440 Direct materials 3,570
Direct labor 5,300 Direct labor 700
Factory overhead ? Factory overhead ?

Determine each of the following for July:

a. Direct and indirect materials used.

Direct materials used and transferred to Work in Process $
Indirect materials used and transferred to Factory Overhead $

b. Direct and indirect labor used.

Direct labor used and transferred to Work in Process $
Indirect labor used and transferred to Factory Overhead $

c. Factory overhead applied (a single overhead rate is used based on direct labor cost).
$

d. Cost of completed Jobs 7-01 and 7-02.
$

e. Assume that in addition to indirect materials and indirect labor, factory overhead of $2,100 was incurred during July. Determine the overapplied or underapplied overhead for July.
$  
Answer #1

Determine each of the following for July:

a. Direct and indirect materials used.

Direct materials used and transferred to Work in Process $37210
Indirect materials used and transferred to Factory Overhead $1200

b. Direct and indirect labor used.

Direct labor used and transferred to Work in Process $13300
Indirect labor used and transferred to Factory Overhead $4300

c. Factory overhead applied (a single overhead rate is used based on direct labor cost).
$2640/4800 = 55%

Factory overhead applied = 13300*55% = 7315

d. Cost of completed Jobs 7-01 and 7-02.
$19840+9675 = 29515

e. Assume that in addition to indirect materials and indirect labor, factory overhead of $2,100 was incurred during July. Determine the overapplied or underapplied overhead for July.
Actual overhead = 1200+4300+2100 = 7600

Under applied overhead = 7600-7315 = 285